A production facility is trying to expand production on a machine that is currently used to produce one part. This item has an annual demand of 1000 units, an annual carrying cost of $10 per unit, and a setup cost of $400. They operate 50 weeks per year and can produce 40 units per week.

What percentage of the time are they using this machine?
Answer #1

Annual Demand = D = 1000/year

Setup Cost = Co = $400

Holding Cost = Cc = $10/year

Economic Production Quantity Q* = √(2DCo/Cc) = √(2*1000*400/10) = 283

Number of Production cycles / year = Annual Demand / Q* = 1000/283 = 3.53 or 4

Hence, units produced per year = 4 * 283 = 1132

Production capacity / year = 50*40 = 2000

% of time the machine is used = units produced per year / production capacity per year = 1132/2000 * 100% = 56.6%